# Are EU low-carbon structural funds efficient in reducing emissions?

### v1

@@ Acknowledgements
Regression by type of polluting gas: trend components
Trend components
| Dependent variable: | ||||||
| CO2pc | CO2i | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.170∗∗ | 0.161∗∗ | 0.002 | 0.172∗∗ | 0.193∗∗∗ | 0.094 |
| (0.078) | (0.066) | (0.062) | (0.073) | (0.062) | (0.058) | |
| Low Carbon Exp. pc | 0.678∗∗∗ | 0.833∗ | 0.234 | 0.926∗∗∗ | 1.337∗∗∗ | 0.617 |
| (0.255) | (0.481) | (0.453) | (0.239) | (0.456) | (0.427) | |
| Low Carbon Exp. pcTER | 0.003 | 0.268∗∗ | 0.013 | 0.084 | 0.097 | 0.110 |
| (0.149) | (0.134) | (0.126) | (0.140) | (0.127) | (0.118) | |
| Low Carbon Exp. pcTER | 1.718∗∗∗ | 0.049 | 2.866∗∗∗ | 2.102∗∗∗ | 0.847 | 2.222∗∗∗ |
| (0.501) | (0.947) | (0.894) | (0.470) | (0.899) | (0.842) | |
| Low Carbon Exp. pcLDR | 0.739∗∗∗ | 0.647∗∗∗ | 0.490∗∗∗ | 0.370∗∗∗ | 0.383∗∗∗ | 0.315∗∗∗ |
| (0.096) | (0.083) | (0.078) | (0.090) | (0.078) | (0.074) | |
| Low Carbon Exp. pcLDR | 2.603∗∗∗ | 3.441∗∗∗ | 2.423∗∗∗ | 1.514∗∗∗ | 2.267∗∗∗ | 1.621∗∗∗ |
| (0.255) | (0.494) | (0.466) | (0.239) | (0.469) | (0.439) | |
| GDP | 0.225∗∗∗ | 0.375∗∗∗ | 0.615∗∗∗ | 0.446∗∗∗ | ||
| (0.019) | (0.022) | (0.018) | (0.021) | |||
| EPS | 0.008 | 0.010∗∗ | 0.004 | 0.009∗ | ||
| (0.005) | (0.005) | (0.005) | (0.005) | |||
| Political Ideology | 0.022∗∗∗ | 0.002 | 0.008∗ | 0.006 | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| GVAA | 0.065∗∗∗ | 0.095∗∗∗ | ||||
| (0.014) | (0.014) | |||||
| GVAB-E | 0.244∗∗∗ | 0.055∗∗∗ | ||||
| (0.017) | (0.016) | |||||
| GVAF | 0.005 | 0.074∗∗∗ | ||||
| (0.016) | (0.015) | |||||
| GVAG-J | 0.032 | 0.143∗∗∗ | ||||
| (0.033) | (0.031) | |||||
| GVAK-N | 0.406∗∗∗ | 0.585∗∗∗ | ||||
| (0.031) | (0.029) | |||||
| GVAO-U | 0.430∗∗∗ | 0.435∗∗∗ | ||||
| (0.037) | (0.035) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.531 | 0.609 | 0.659 | 0.760 | 0.763 | 0.796 |
| Dependent variable: | ||||||
| CH4pc | CH4i | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.134∗ | 0.100 | 0.103 | 0.136∗ | 0.132∗∗ | 0.199∗∗∗ |
| (0.079) | (0.064) | (0.064) | (0.075) | (0.065) | (0.066) | |
| Low Carbon Exp. pc | 0.378 | 0.077 | 0.114 | 0.130 | 0.580 | 0.965∗∗ |
| (0.258) | (0.471) | (0.468) | (0.247) | (0.473) | (0.483) | |
| Low Carbon Exp. pcTER | 0.037 | 0.033 | 0.044 | 0.044 | 0.138 | 0.080 |
| (0.151) | (0.131) | (0.130) | (0.145) | (0.132) | (0.134) | |
| Low Carbon Exp. pcTER | 2.751∗∗∗ | 2.460∗∗∗ | 3.132∗∗∗ | 3.133∗∗∗ | 3.258∗∗∗ | 2.488∗∗∗ |
| (0.507) | (0.928) | (0.923) | (0.487) | (0.933) | (0.952) | |
| Low Carbon Exp. pcLDR | 0.406∗∗∗ | 0.429∗∗∗ | 0.352∗∗∗ | 0.038 | 0.165∗∗ | 0.177∗∗ |
| (0.097) | (0.081) | (0.081) | (0.093) | (0.081) | (0.083) | |
| Low Carbon Exp. pcLDR | 0.404 | 1.339∗∗∗ | 0.836∗ | 0.683∗∗∗ | 0.163 | 0.034 |
| (0.258) | (0.484) | (0.481) | (0.248) | (0.486) | (0.496) | |
| GDP | 0.172∗∗∗ | 0.100∗∗∗ | 0.668∗∗∗ | 0.722∗∗∗ | ||
| (0.020) | (0.021) | (0.019) | (0.021) | |||
| EPS | 0.026∗∗∗ | 0.019∗∗∗ | 0.030∗∗∗ | 0.021∗∗∗ | ||
| (0.005) | (0.005) | (0.005) | (0.005) | |||
| Political Ideology | 0.003 | 0.003 | 0.017∗∗∗ | 0.011∗∗ | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| GVAA | 0.023 | 0.007 | ||||
| (0.015) | (0.015) | |||||
| GVAB-E | 0.108∗∗∗ | 0.191∗∗∗ | ||||
| (0.018) | (0.018) | |||||
| GVAF | 0.136∗∗∗ | 0.214∗∗∗ | ||||
| (0.016) | (0.017) | |||||
| GVAG-J | 0.112∗∗∗ | 0.001 | ||||
| (0.034) | (0.036) | |||||
| GVAK-N | 0.010 | 0.189∗∗∗ | ||||
| (0.032) | (0.033) | |||||
| GVAO-U | 0.170∗∗∗ | 0.175∗∗∗ | ||||
| (0.038) | (0.039) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.345 | 0.473 | 0.490 | 0.683 | 0.716 | 0.711 |
| Dependent variable: | ||||||
| N2Opc | N2Oi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.108 | 0.060 | 0.013 | 0.102∗ | 0.082 | 0.111∗∗ |
| (0.071) | (0.055) | (0.054) | (0.061) | (0.053) | (0.055) | |
| Low Carbon Exp. pc | 0.189 | 0.052 | 0.553 | 0.366∗ | 0.452 | 0.291 |
| (0.233) | (0.403) | (0.397) | (0.199) | (0.386) | (0.402) | |
| Low Carbon Exp. pcTER | 0.254∗ | 0.036 | 0.104 | 0.353∗∗∗ | 0.240∗∗ | 0.021 |
| (0.136) | (0.113) | (0.110) | (0.116) | (0.108) | (0.111) | |
| Low Carbon Exp. pcTER | 1.804∗∗∗ | 1.333∗ | 2.032∗∗∗ | 2.275∗∗∗ | 2.352∗∗∗ | 1.364∗ |
| (0.457) | (0.795) | (0.781) | (0.391) | (0.760) | (0.792) | |
| Low Carbon Exp. pcLDR | 0.693∗∗∗ | 0.344∗∗∗ | 0.290∗∗∗ | 0.331∗∗∗ | 0.085 | 0.116∗ |
| (0.087) | (0.069) | (0.068) | (0.075) | (0.066) | (0.069) | |
| Low Carbon Exp. pcLDR | 2.399∗∗∗ | 1.722∗∗∗ | 1.273∗∗∗ | 1.297∗∗∗ | 0.534 | 0.465 |
| (0.233) | (0.414) | (0.407) | (0.199) | (0.396) | (0.412) | |
| GDP | 0.103∗∗∗ | 0.135∗∗∗ | 0.759∗∗∗ | 0.714∗∗∗ | ||
| (0.018) | (0.019) | (0.015) | (0.018) | |||
| EPS | 0.006 | 0.010∗∗ | 0.010∗∗ | 0.012∗∗ | ||
| (0.004) | (0.004) | (0.004) | (0.005) | |||
| Political Ideology | 0.016∗∗∗ | 0.010∗∗∗ | 0.003 | 0.001 | ||
| (0.004) | (0.004) | (0.003) | (0.004) | |||
| GVAA | 0.091∗∗∗ | 0.062∗∗∗ | ||||
| (0.013) | (0.013) | |||||
| GVAB-E | 0.113∗∗∗ | 0.186∗∗∗ | ||||
| (0.015) | (0.015) | |||||
| GVAF | 0.040∗∗∗ | 0.039∗∗∗ | ||||
| (0.014) | (0.014) | |||||
| GVAG-J | 0.009 | 0.101∗∗∗ | ||||
| (0.029) | (0.030) | |||||
| GVAK-N | 0.188∗∗∗ | 0.369∗∗∗ | ||||
| (0.027) | (0.027) | |||||
| GVAO-U | 0.032 | 0.042 | ||||
| (0.036) | (0.036) | |||||
| Observations | 3,856 | 2,982 | 2,982 | 3,856 | 2,982 | 2,982 |
| Adjusted R2 | 0.109 | 0.159 | 0.203 | 0.594 | 0.619 | 0.595 |
Cycle components
| Dependent variable: | ||||||
| CO2pc | CO2i | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.072 | 0.003 | 0.008 | 0.059 | 0.001 | 0.004 |
| (0.054) | (0.054) | (0.054) | (0.054) | (0.054) | (0.054) | |
| Low Carbon Exp. pc | 0.162 | 0.482 | 0.394 | 0.152 | 0.463 | 0.315 |
| (0.177) | (0.395) | (0.394) | (0.177) | (0.396) | (0.397) | |
| Low Carbon Exp. pcTER | 0.096 | 0.308∗∗∗ | 0.374∗∗∗ | 0.089 | 0.311∗∗∗ | 0.377∗∗∗ |
| (0.104) | (0.108) | (0.108) | (0.103) | (0.108) | (0.108) | |
| Low Carbon Exp. pcTER | 0.115 | 1.745∗∗ | 1.918∗∗ | 0.134 | 1.871∗∗ | 1.668∗∗ |
| (0.344) | (0.763) | (0.760) | (0.342) | (0.763) | (0.765) | |
| Low Carbon Exp. pcLDR | 0.062 | 0.052 | 0.063 | 0.069 | 0.067 | 0.067 |
| (0.067) | (0.068) | (0.068) | (0.067) | (0.068) | (0.068) | |
| Low Carbon Exp. pcLDR | 0.024 | 0.681∗ | 0.721∗ | 0.030 | 0.674∗ | 0.494 |
| (0.179) | (0.407) | (0.407) | (0.178) | (0.408) | (0.410) | |
| GDP | 0.567∗∗∗ | 0.660∗∗∗ | 0.428∗∗∗ | 0.323∗∗∗ | ||
| (0.034) | (0.042) | (0.034) | (0.042) | |||
| EPS | 0.036∗∗∗ | 0.041∗∗∗ | 0.028∗∗∗ | 0.028∗∗∗ | ||
| (0.004) | (0.005) | (0.004) | (0.005) | |||
| Political Ideology | 0.001 | 0.0004 | 0.003 | 0.002 | ||
| (0.003) | (0.003) | (0.003) | (0.003) | |||
| GVAA | 0.006 | 0.010 | ||||
| (0.010) | (0.010) | |||||
| GVAB-E | 0.125∗∗∗ | 0.069∗∗∗ | ||||
| (0.019) | (0.019) | |||||
| GVAF | 0.100∗∗∗ | 0.044∗∗∗ | ||||
| (0.015) | (0.015) | |||||
| GVAG-J | 0.078∗∗∗ | 0.179∗∗∗ | ||||
| (0.030) | (0.030) | |||||
| GVAK-N | 0.125∗∗∗ | 0.087∗∗ | ||||
| (0.039) | (0.039) | |||||
| GVAO-U | 0.291∗∗∗ | 0.103∗∗ | ||||
| (0.044) | (0.044) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.206 | 0.261 | 0.270 | 0.044 | 0.042 | 0.041 |
| Dependent variable: | ||||||
| CH4pc | CH4i | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.198∗∗∗ | 0.174∗∗∗ | 0.184∗∗∗ | 0.185∗∗∗ | 0.173∗∗∗ | 0.172∗∗∗ |
| (0.047) | (0.048) | (0.048) | (0.047) | (0.047) | (0.048) | |
| Low Carbon Exp. pc | 0.075 | 0.188 | 0.168 | 0.085 | 0.207 | 0.247 |
| (0.152) | (0.348) | (0.348) | (0.152) | (0.346) | (0.353) | |
| Low Carbon Exp. pcTER | 0.152∗ | 0.237∗∗ | 0.256∗∗∗ | 0.146 | 0.240∗∗ | 0.259∗∗∗ |
| (0.089) | (0.095) | (0.095) | (0.089) | (0.094) | (0.096) | |
| Low Carbon Exp. pcTER | 0.699∗∗ | 1.353∗∗ | 1.455∗∗ | 0.718∗∗ | 1.479∗∗ | 1.206∗ |
| (0.295) | (0.671) | (0.670) | (0.294) | (0.667) | (0.681) | |
| Low Carbon Exp. pcLDR | 0.094 | 0.066 | 0.069 | 0.086 | 0.052 | 0.066 |
| (0.058) | (0.060) | (0.060) | (0.057) | (0.060) | (0.061) | |
| Low Carbon Exp. pcLDR | 0.199 | 0.198 | 0.124 | 0.194 | 0.190 | 0.105 |
| (0.153) | (0.359) | (0.359) | (0.153) | (0.356) | (0.364) | |
| GDP | 0.163∗∗∗ | 0.181∗∗∗ | 0.831∗∗∗ | 0.802∗∗∗ | ||
| (0.029) | (0.037) | (0.029) | (0.037) | |||
| EPS | 0.004 | 0.008∗∗ | 0.004 | 0.006 | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| Political Ideology | 0.002 | 0.001 | 0.002 | 0.001 | ||
| (0.003) | (0.003) | (0.003) | (0.003) | |||
| GVAA | 0.009 | 0.007 | ||||
| (0.009) | (0.009) | |||||
| GVAB-E | 0.048∗∗∗ | 0.146∗∗∗ | ||||
| (0.016) | (0.017) | |||||
| GVAF | 0.074∗∗∗ | 0.018 | ||||
| (0.013) | (0.013) | |||||
| GVAG-J | 0.002 | 0.260∗∗∗ | ||||
| (0.026) | (0.027) | |||||
| GVAK-N | 0.011 | 0.223∗∗∗ | ||||
| (0.034) | (0.035) | |||||
| GVAO-U | 0.045 | 0.233∗∗∗ | ||||
| (0.039) | (0.039) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.034 | 0.040 | 0.032 | 0.258 | 0.243 | 0.216 |
| Dependent variable: | ||||||
| N2Opc | N2Oi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.059 | 0.035 | 0.062 | 0.045 | 0.032 | 0.049 |
| (0.056) | (0.063) | (0.063) | (0.055) | (0.063) | (0.064) | |
| Low Carbon Exp. pc | 0.050 | 0.160 | 0.157 | 0.062 | 0.145 | 0.076 |
| (0.181) | (0.462) | (0.461) | (0.180) | (0.458) | (0.465) | |
| Low Carbon Exp. pcTER | 0.072 | 0.076 | 0.101 | 0.080 | 0.074 | 0.098 |
| (0.106) | (0.126) | (0.126) | (0.105) | (0.125) | (0.127) | |
| Low Carbon Exp. pcTER | 0.245 | 0.718 | 1.139 | 0.263 | 0.844 | 0.893 |
| (0.351) | (0.891) | (0.889) | (0.348) | (0.883) | (0.896) | |
| Low Carbon Exp. pcLDR | 0.191∗∗∗ | 0.214∗∗∗ | 0.193∗∗ | 0.184∗∗∗ | 0.200∗∗ | 0.190∗∗ |
| (0.069) | (0.080) | (0.080) | (0.068) | (0.079) | (0.080) | |
| Low Carbon Exp. pcLDR | 0.065 | 0.050 | 0.016 | 0.070 | 0.057 | 0.250 |
| (0.183) | (0.476) | (0.476) | (0.181) | (0.472) | (0.480) | |
| GDP | 0.140∗∗∗ | 0.157∗∗∗ | 0.856∗∗∗ | 0.826∗∗∗ | ||
| (0.035) | (0.049) | (0.035) | (0.049) | |||
| EPS | 0.0002 | 0.004 | 0.007 | 0.017∗∗∗ | ||
| (0.005) | (0.005) | (0.005) | (0.005) | |||
| Political Ideology | 0.009∗∗ | 0.007∗ | 0.005 | 0.005 | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| GVAA | 0.029∗∗ | 0.012 | ||||
| (0.012) | (0.012) | |||||
| GVAB-E | 0.039∗ | 0.156∗∗∗ | ||||
| (0.022) | (0.022) | |||||
| GVAF | 0.019 | 0.038∗∗ | ||||
| (0.017) | (0.018) | |||||
| GVAG-J | 0.114∗∗∗ | 0.147∗∗∗ | ||||
| (0.035) | (0.035) | |||||
| GVAK-N | 0.166∗∗∗ | 0.383∗∗∗ | ||||
| (0.046) | (0.046) | |||||
| GVAO-U | 0.031 | 0.154∗∗∗ | ||||
| (0.052) | (0.052) | |||||
| Observations | 3,856 | 2,982 | 2,982 | 3,856 | 2,982 | 2,982 |
| Adjusted R2 | 0.018 | 0.011 | 0.020 | 0.241 | 0.175 | 0.156 |

### v2

@@ Acknowledgements
The cohesion data platform presents estimations of the expenditure by NUTS 2, which are aggregated for the most significant types of structural funds only. Therefore, to address our need for yearly data, we have opted for two approaches: the one presented in the main text, which assumes that expenditures are distributed evenly over time (dividing the reported expenditures over the programming period), and the another one, presented here, which assumes they follow the same pattern as the total aggregated expenditure reported by the cohesion data platform.888Data available at: https://cohesiondata.ec.europa.eu/Other/Historic-EU-payments-regionalised-and-modelled/tc55-7ysv/about_data. Both strategies yielded qualitatively similar conclusions, as a robustness check the results of the second strategy are shown in Tables A2 and A3.
@@
Testing for instruments exogeneity
| Dependent variable: | ||||
| Trend residual | Cycle residual | |||
| (3t) | (6t) | (3c) | (6c) | |
| GHGEpc | GHGEi | GHGEpc | GHGEi | |
| Labour productivity ppet-8 | 0.003 | 0.002 | -0.002 | 0.000 |
| (0.004) | (0.004) | (0.003) | (0.003) | |
| Consumption of fixed capitalt-8 | -0.001 | -0.002 | 0.004 | 0.001 |
| (0.005) | (0.005) | (0.004) | (0.004) | |
| Hours worked per capitat-8 | 0.000 | 0.000 | -0.000 | -0.000 |
| (0.000) | (0.000) | (0.000) | (0.000) | |
| Compensation of Employeest-8 | 0.000 | 0.002 | -0.004 | -0.001 |
| (0.005) | (0.005) | (0.004) | (0.004) | |
| TER | 0.002 | 0.003 | -0.003 | 0.000 |
| (0.004) | (0.004) | (0.003) | (0.003) | |
| LDR | 0.000 | 0.002 | -0.002 | -0.000 |
| (0.003) | (0.003) | (0.002) | (0.002) | |
| Constant | -0.031 | -0.039 | 0.036 | -0.001 |
| (0.043) | (0.044) | (0.032) | (0.032) | |
| Observations | 3,408 | 3,408 | 3,408 | 3,408 |
| R-squared | 0.000 | 0.001 | 0.001 | 0.000 |
| Notes: Standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||